$SPAD/input schaum30.input

Timothy Daly December 3, 2016

Contents

2

1 [?]:14.604 ∫ tanh ax dx

∫ tanh ax = \frac{1}{a} ln cosh ax

— * —

)set break resume
)spool schaum30.output
)set message test on
)set message auto off
)clear all
--S 1 of 46
aa:=integrate(tanh(a*x),x)
--R
--R
--R                    2cosh(a x)
--R        log(- ---------------------) - a x
--R              sinh(a x) - cosh(a x)
--R   (1)  ----------------------------------
--R                         a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 2 of 46
bb:=1/a*log(cosh(a*x))
--R
--R        log(cosh(a x))
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E
--S 3 of 46
cc:=aa-bb
--R
--R                                       2cosh(a x)
--R        - log(cosh(a x)) + log(- ---------------------) - a x
--R                                 sinh(a x) - cosh(a x)
--R   (3)  -----------------------------------------------------
--R                                  a
--R                                                    Type: Expression(Integer)
--E
--S 4 of 46
dd:=expandLog cc
--R
--R        - log(sinh(a x) - cosh(a x)) + log(- 2) - a x

--R   (4)  ---------------------------------------------
--R                              a
--R                                                    Type: Expression(Integer)
--E
--S 5 of 46      14:604 Schaums and Axiom differ by a constant
ee:=complexNormalize dd
--R
--R        - log(- 1) + log(- 2)
--R   (5)  ---------------------
--R                  a
--R                                                    Type: Expression(Integer)
--E

———-

2 [?]:14.605 ∫ tanh^{2} ax dx

∫ tanh^{2} ax = x − \frac{tanh ax}{a}

— * —

)clear all
--S 6 of 46
aa:=integrate(tanh(a*x)^2,x)
--R
--R
--R        - sinh(a x) + (a x + 1)cosh(a x)
--R   (1)  --------------------------------
--R                   a cosh(a x)
--R                                         Type: Union(Expression(Integer),...)
--E
--S 7 of 46
bb:=x-tanh(a*x)/a
--R
--R        - tanh(a x) + a x
--R   (2)  -----------------
--R                a
--R                                                    Type: Expression(Integer)
--E
--S 8 of 46
cc:=aa-bb
--R
--R        cosh(a x)tanh(a x) - sinh(a x) + cosh(a x)
--R   (3)  ------------------------------------------

--R                        a cosh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 9 of 46
tanhrule:=rule(tanh(x) == sinh(x)/cosh(x))
--R
--R                   sinh(x)
--R   (4)  tanh(x) == -------
--R                   cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 10 of 46     14:605 Schaums and Axiom differ by a constant
dd:=tanhrule cc
--R
--R        1
--R   (5)  -
--R        a
--R                                                    Type: Expression(Integer)
--E

———-

3 [?]:14.606 ∫ tanh^{3} ax dx

∫ tanh^{3} ax = \frac{1}{a} ln cosh ax − \frac{tanh^{2} ax}{2a}

— * —

)clear all
--S 11 of 46
aa:=integrate(tanh(a*x)^3,x)
--R
--R
--R   (1)
--R                    4                      3              2              2
--R           sinh(a x)  + 4cosh(a x)sinh(a x)  + (6cosh(a x)  + 2)sinh(a x)
--R         +
--R                      3                                   4             2
--R           (4cosh(a x)  + 4cosh(a x))sinh(a x) + cosh(a x)  + 2cosh(a x)  + 1
--R      *
--R                     2cosh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R                      4                          3

--R       - a x sinh(a x)  - 4a x cosh(a x)sinh(a x)
--R     +
--R                        2                     2
--R       (- 6a x cosh(a x)  - 2a x + 2)sinh(a x)
--R     +
--R                        3                                                  4
--R       (- 4a x cosh(a x)  + (- 4a x + 4)cosh(a x))sinh(a x) - a x cosh(a x)
--R     +
--R                            2
--R       (- 2a x + 2)cosh(a x)  - a x
--R  /
--R                  4                        3                2               2
--R       a sinh(a x)  + 4a cosh(a x)sinh(a x)  + (6a cosh(a x)  + 2a)sinh(a x)
--R     +
--R                  3                                       4               2
--R     (4a cosh(a x)  + 4a cosh(a x))sinh(a x) + a cosh(a x)  + 2a cosh(a x)  + a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 12 of 46
bb:=1/a*log(cosh(a*x))-tanh(a*x)^2/(2*a)
--R
--R                                   2
--R        2log(cosh(a x)) - tanh(a x)
--R   (2)  ----------------------------
--R                     2a
--R                                                    Type: Expression(Integer)
--E
--S 13 of 46     14:606 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                       4                      3                 2              2
--R           - 2sinh(a x)  - 8cosh(a x)sinh(a x)  + (- 12cosh(a x)  - 4)sinh(a x)
--R         +
--R                      3                                    4             2
--R         (- 8cosh(a x)  - 8cosh(a x))sinh(a x) - 2cosh(a x)  - 4cosh(a x)  - 2
--R      *
--R         log(cosh(a x))
--R     +
--R                     4                      3               2              2
--R           2sinh(a x)  + 8cosh(a x)sinh(a x)  + (12cosh(a x)  + 4)sinh(a x)
--R         +
--R                      3                                    4             2
--R           (8cosh(a x)  + 8cosh(a x))sinh(a x) + 2cosh(a x)  + 4cosh(a x)  + 2
--R      *
--R                     2cosh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)

--R     +
--R                    4                      3              2              2
--R           sinh(a x)  + 4cosh(a x)sinh(a x)  + (6cosh(a x)  + 2)sinh(a x)
--R         +
--R                      3                                   4             2
--R           (4cosh(a x)  + 4cosh(a x))sinh(a x) + cosh(a x)  + 2cosh(a x)  + 1
--R      *
--R                  2
--R         tanh(a x)
--R     +
--R                       4                          3
--R       - 2a x sinh(a x)  - 8a x cosh(a x)sinh(a x)
--R     +
--R                         2                     2
--R       (- 12a x cosh(a x)  - 4a x + 4)sinh(a x)
--R     +
--R                        3                                                   4
--R       (- 8a x cosh(a x)  + (- 8a x + 8)cosh(a x))sinh(a x) - 2a x cosh(a x)
--R     +
--R                            2
--R       (- 4a x + 4)cosh(a x)  - 2a x
--R  /
--R                   4                        3                 2               2
--R       2a sinh(a x)  + 8a cosh(a x)sinh(a x)  + (12a cosh(a x)  + 4a)sinh(a x)
--R     +
--R                    3                                        4               2
--R       (8a cosh(a x)  + 8a cosh(a x))sinh(a x) + 2a cosh(a x)  + 4a cosh(a x)
--R     +
--R       2a
--R                                                    Type: Expression(Integer)
--E

———-

4 [?]:14.607 ∫ tanh^{n} ax sech^{2}ax dx

∫ tanh^{n} ax sech^{2}ax = \frac{tanh^{n+1} ax}{(n + 1)a}

— * —

)clear all
--S 14 of 46
aa:=integrate(tanh(a*x)^n*sech(a*x)^2,x)
--R
--R
--R                            sinh(a x)                         sinh(a x)
--R        sinh(a x)sinh(n log(---------)) + sinh(a x)cosh(n log(---------))

--R                            cosh(a x)                         cosh(a x)
--R   (1)  -----------------------------------------------------------------
--R                                (a n + a)cosh(a x)
--R                                         Type: Union(Expression(Integer),...)
--E
--S 15 of 46
bb:=tanh(a*x)^(n+1)/((n+1)*a)
--R
--R                 n + 1
--R        tanh(a x)
--R   (2)  --------------
--R            a n + a
--R                                                    Type: Expression(Integer)
--E
--S 16 of 46     14:607 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                           sinh(a x)                         sinh(a x)
--R       sinh(a x)sinh(n log(---------)) + sinh(a x)cosh(n log(---------))
--R                           cosh(a x)                         cosh(a x)
--R     +
--R                           n + 1
--R       - cosh(a x)tanh(a x)
--R  /
--R     (a n + a)cosh(a x)
--R                                                    Type: Expression(Integer)
--E

———-

5 [?]:14.608 ∫ \begin{matrix}sech & 2 & ax \\ tanh & ax \end{matrix} dx

∫ \begin{matrix}sech & 2 & ax \\ tanh & ax \end{matrix} = \frac{1}{a} ln tanh ax

— * —

)clear all
--S 17 of 46
aa:=integrate(sech(a*x)^2/tanh(a*x),x)
--R
--R
--R                      2cosh(a x)                     2sinh(a x)
--R        - log(- ---------------------) + log(- ---------------------)

--R                sinh(a x) - cosh(a x)          sinh(a x) - cosh(a x)
--R   (1)  -------------------------------------------------------------
--R                                      a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 18 of 46
bb:=1/a*log(tanh(a*x))
--R
--R        log(tanh(a x))
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E
--S 19 of 46
cc:=aa-bb
--R
--R   (3)
--R                                      2cosh(a x)
--R       - log(tanh(a x)) - log(- ---------------------)
--R                                sinh(a x) - cosh(a x)
--R     +
--R                   2sinh(a x)
--R       log(- ---------------------)
--R             sinh(a x) - cosh(a x)
--R  /
--R     a
--R                                                    Type: Expression(Integer)
--E
--S 20 of 46
tanhrule:=rule(tanh(x) == sinh(x)/cosh(x))
--R
--R                   sinh(x)
--R   (4)  tanh(x) == -------
--R                   cosh(x)
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 21 of 46
dd:=tanhrule cc
--R
--R   (5)
--R             sinh(a x)                2cosh(a x)
--R       - log(---------) - log(- ---------------------)
--R             cosh(a x)          sinh(a x) - cosh(a x)
--R     +
--R                   2sinh(a x)
--R       log(- ---------------------)

--R             sinh(a x) - cosh(a x)
--R  /
--R     a
--R                                                    Type: Expression(Integer)
--E
--S 22 of 46     14:608 Schaums and Axiom agree
ee:=expandLog dd
--R
--R   (6)  0
--R                                                    Type: Expression(Integer)
--E

———-

6 [?]:14.609 ∫ \frac{dx}{tanh ax} dx

∫ \frac{1}{tanh ax} = \frac{1}{a} ln sinh ax

— * —

)clear all
--S 23 of 46
aa:=integrate(1/tanh(a*x),x)
--R
--R
--R                    2sinh(a x)
--R        log(- ---------------------) - a x
--R              sinh(a x) - cosh(a x)
--R   (1)  ----------------------------------
--R                         a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 24 of 46
bb:=1/a*log(sinh(a*x))
--R
--R        log(sinh(a x))
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E
--S 25 of 46
cc:=aa-bb
--R

--R                                       2sinh(a x)
--R        - log(sinh(a x)) + log(- ---------------------) - a x
--R                                 sinh(a x) - cosh(a x)
--R   (3)  -----------------------------------------------------
--R                                  a
--R                                                    Type: Expression(Integer)
--E
--S 26 of 46
dd:=expandLog cc
--R
--R        - log(sinh(a x) - cosh(a x)) + log(- 2) - a x
--R   (4)  ---------------------------------------------
--R                              a
--R                                                    Type: Expression(Integer)
--E
--S 27 of 46     14:609 Schaums and Axiom differ by a constant
ee:=complexNormalize dd
--R
--R        - log(- 1) + log(- 2)
--R   (5)  ---------------------
--R                  a
--R                                                    Type: Expression(Integer)
--E

———-

7 [?]:14.610 ∫ x tanh ax dx

∫ x tanh ax = \frac{1}{a^{2}} \{\frac{(ax)^{3}}{3} − \frac{(ax)^{5}}{15} + \frac{2(ax)^{7}}{105} − ⋯ \frac{(−1)^{n−1}2^{2n}(2^{2n} − 1)B_{n}(ax)^{2n+1}}{(2n + 1)!} + ⋯\}

— * —

)clear all
--S 28 of 46     14:610 Axiom cannot compute this integral
aa:=integrate(x*tanh(a*x),x)
--R
--R
--R           x
--R         ++
--I   (1)   |   %O tanh(%O a)d%O
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E

———-

8 [?]:14.611 ∫ x tanh^{2} ax dx

∫ x tanh^{2} ax = \frac{x^{2}}{2} − \frac{x tanh ax}{a} + \frac{1}{a^{2}} ln cosh ax

— * —

)clear all
--S 29 of 46
aa:=integrate(x*tanh(a*x)^2,x)
--R
--R
--R   (1)
--R                    2                                   2
--R         (2sinh(a x)  + 4cosh(a x)sinh(a x) + 2cosh(a x)  + 2)
--R      *
--R                     2cosh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R         2 2                 2      2 2
--R       (a x  - 4a x)sinh(a x)  + (2a x  - 8a x)cosh(a x)sinh(a x)
--R     +
--R         2 2                 2    2 2
--R       (a x  - 4a x)cosh(a x)  + a x
--R  /
--R       2         2     2                       2         2     2
--R     2a sinh(a x)  + 4a cosh(a x)sinh(a x) + 2a cosh(a x)  + 2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 30 of 46
bb:=x^2/2-(x*tanh(a*x))/a+1/a^2*log(cosh(a*x))
--R
--R                                            2 2
--R        2log(cosh(a x)) - 2a x tanh(a x) + a x
--R   (2)  ---------------------------------------
--R                            2
--R                          2a
--R                                                    Type: Expression(Integer)
--E
--S 31 of 46
cc:=aa-bb
--R
--R   (3)
--R                   2                                  2
--R       (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  - 1)log(cosh(a x))

--R     +
--R                   2                                  2
--R         (sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  + 1)
--R      *
--R                     2cosh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R                       2                                          2
--R         (a x sinh(a x)  + 2a x cosh(a x)sinh(a x) + a x cosh(a x)  + a x)
--R      *
--R         tanh(a x)
--R     +
--R                       2                                           2
--R       - 2a x sinh(a x)  - 4a x cosh(a x)sinh(a x) - 2a x cosh(a x)
--R  /
--R      2         2     2                      2         2    2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  + a
--R                                                    Type: Expression(Integer)
--E
--S 32 of 46
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (4)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 33 of 46
dd:=sinhsqrrule cc
--R
--R   (5)
--R                                                       2
--R       (- 4cosh(a x)sinh(a x) - cosh(2a x) - 2cosh(a x)  - 1)log(cosh(a x))
--R     +
--R                                                       2
--R         (4cosh(a x)sinh(a x) + cosh(2a x) + 2cosh(a x)  + 1)
--R      *
--R                     2cosh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R                                                                   2
--R         (4a x cosh(a x)sinh(a x) + a x cosh(2a x) + 2a x cosh(a x)  + a x)
--R      *
--R         tanh(a x)
--R     +
--R                                                                   2

--R       - 8a x cosh(a x)sinh(a x) - 2a x cosh(2a x) - 4a x cosh(a x)  + 2a x
--R  /
--R       2                      2               2         2    2
--R     4a cosh(a x)sinh(a x) + a cosh(2a x) + 2a cosh(a x)  + a
--R                                                    Type: Expression(Integer)
--E
--S 34 of 46
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (6)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 35 of 46
ee:=coshsqrrule dd
--R
--R   (7)
--R       (- 2cosh(a x)sinh(a x) - cosh(2a x) - 1)log(cosh(a x))
--R     +
--R                                                         2cosh(a x)
--R       (2cosh(a x)sinh(a x) + cosh(2a x) + 1)log(- ---------------------)
--R                                                   sinh(a x) - cosh(a x)
--R     +
--R       (2a x cosh(a x)sinh(a x) + a x cosh(2a x) + a x)tanh(a x)
--R     +
--R       - 4a x cosh(a x)sinh(a x) - 2a x cosh(2a x)
--R  /
--R       2                      2              2
--R     2a cosh(a x)sinh(a x) + a cosh(2a x) + a
--R                                                    Type: Expression(Integer)
--E
--S 36 of 46
ff:=expandLog ee
--R
--R   (8)
--R       (- 2cosh(a x)sinh(a x) - cosh(2a x) - 1)log(sinh(a x) - cosh(a x))
--R     +
--R       (2a x cosh(a x)sinh(a x) + a x cosh(2a x) + a x)tanh(a x)
--R     +
--R       (2log(- 2) - 4a x)cosh(a x)sinh(a x) + (log(- 2) - 2a x)cosh(2a x)
--R     +
--R       log(- 2)
--R  /
--R       2                      2              2
--R     2a cosh(a x)sinh(a x) + a cosh(2a x) + a
--R                                                    Type: Expression(Integer)

--E
--S 37 of 46
sinhcoshrule:=rule(sinh(x)*cosh(y) == 1/2*(sinh(x+y)+sinh(x-y)))
--R
--I                             %N sinh(y + x) - %N sinh(y - x)
--I   (9)  %N cosh(y)sinh(x) == -------------------------------
--R                                            2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 38 of 46
gg:=sinhcoshrule ff
--R
--R   (10)
--R       (- sinh(2a x) - cosh(2a x) - 1)log(sinh(a x) - cosh(a x))
--R     +
--R       (a x sinh(2a x) + a x cosh(2a x) + a x)tanh(a x)
--R     +
--R       (log(- 2) - 2a x)sinh(2a x) + (log(- 2) - 2a x)cosh(2a x) + log(- 2)
--R  /
--R      2              2              2
--R     a sinh(2a x) + a cosh(2a x) + a
--R                                                    Type: Expression(Integer)
--E
--S 39 of 46     14:611 Schaums and Axiom differ by a constant
hh:=complexNormalize gg
--R
--R         - log(- 1) + log(- 2)
--R   (11)  ---------------------
--R                    2
--R                   a
--R                                                    Type: Expression(Integer)
--E

———-

9 [?]:14.612 ∫ \frac{tanh ax}{x} dx

∫ \frac{tanh ax}{x} = ax − \frac{(ax)^{3}}{9} + \frac{2(ax)^{5}}{75} − ⋯ \frac{(−1)^{n−1}2^{2n}(2^{2n} − 1)B_{n}(ax)^{2n−1}}{(2n − 1)(2n)!} + ⋯

— * —

)clear all
--S 40 of 46     14:612 Axiom cannot compute this integral

aa:=integrate(tanh(a*x)/x,x)
--R
--R
--R           x
--I         ++  tanh(%O a)
--I   (1)   |   ---------- d%O
--I        ++       %O
--R                                         Type: Union(Expression(Integer),...)
--E

———-

10 [?]:14.613 ∫ \frac{dx}{p + q tanh ax} dx

∫ \frac{1}{p + q tanh ax} = \frac{px}{p^{2} − q^{2}} − \frac{q}{a(p^{2} − q^{2})} ln(q sinh ax + p cosh ax)

— * —

)clear all
--S 41 of 46
aa:=integrate(1/(p+q*tanh(a*x)),x)
--R
--R
--R              - 2q sinh(a x) - 2p cosh(a x)
--R        q log(-----------------------------) + (- a q - a p)x
--R                  sinh(a x) - cosh(a x)
--R   (1)  -----------------------------------------------------
--R                                2      2
--R                             a q  - a p
--R                                         Type: Union(Expression(Integer),...)
--E
--S 42 of 46
bb:=(p*x)/(p^2-q^2)-q/(a*(p^2-q^2))*log(q*sinh(a*x)+p*cosh(a*x))
--R
--R        q log(q sinh(a x) + p cosh(a x)) - a p x
--R   (2)  ----------------------------------------
--R                          2      2
--R                       a q  - a p
--R                                                    Type: Expression(Integer)
--E
--S 43 of 46
cc:=aa-bb
--R
--R   (3)

--R                                                  - 2q sinh(a x) - 2p cosh(a x)
--R       - q log(q sinh(a x) + p cosh(a x)) + q log(-----------------------------)
--R                                                      sinh(a x) - cosh(a x)
--R     +
--R       - a q x
--R  /
--R        2      2
--R     a q  - a p
--R                                                    Type: Expression(Integer)
--E
--S 44 of 46
dd:=expandLog cc
--R
--R   (4)
--R       - q log(q sinh(a x) + p cosh(a x)) - q log(sinh(a x) - cosh(a x))
--R     +
--R       q log(- q sinh(a x) - p cosh(a x)) + q log(2) - a q x
--R  /
--R        2      2
--R     a q  - a p
--R                                                    Type: Expression(Integer)
--E
--S 45 of 46     14:613 Schaums and Axiom differ by a constant
ee:=complexNormalize dd
--R
--R        q log(2) - 2q log(- 1)
--R   (5)  ----------------------
--R                 2      2
--R              a q  - a p
--R                                                    Type: Expression(Integer)
--E

———-

11 [?]:14.614 ∫ tanh^{n} ax dx

∫ tanh^{n} ax = \frac{− tanh^{n−1} ax}{a(n − 1)} + ∫ tanh^{n−2} ax

— * —

)clear all
--S 46 of 46     14:614 Axiom cannot compute this integral
aa:=integrate(tanh(a*x)^n,x)
--R
--R

--R           x
--R         ++            n
--I   (1)   |   tanh(%O a) d%O
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E
)spool
)lisp (bye)

———-

References

[1] Spiegel, Murray R. Mathematical Handbook of Formulas and Tables Schaum’s Outline Series McGraw-Hill 1968 pp89-90